OpenStudy (pandaecat):

Which of the following is equal to the square root of the cube root of 6?

OpenStudy (pandaecat):

\[6\frac{ 1 }{ 6 }\] \[6\frac{ 1 }{ 3 }\] \[6\frac{ 2 }{ 3}\] \[6\frac{ 3 }{ 2}\]